Which of the following aqueous solutions when mixed will have a precipitation reaction occur? If yes write balanced molecular, total ionic and net ionic equations.
a) Potassium carbonate + barium hydroxide
b) aluminum nitrate + sodium phosphate
1
Expert's answer
2020-09-29T06:17:13-0400

Precipitate is formed in both reactions:

a) Molecular equation:

K2CO3+Ba(OH)2=2KOH+BaCO3

Total ionic equation:

2K+ + CO32- +Ba2+ + 2OH- = 2K+ + 2OH- + BaCO3

Net ionic equation:

CO32- + Ba2+ = BaCO3.

b) Molecular equation:

Al(NO3)3+Na3PO4 = AlPO4 + 3NaNO3

Total ionic equation:

Al3+ + 3NO3- + 3Na+ + PO43- = AlPO4 + 3Na+ + 3NO3-

Net ionic equation:

Al3+ + PO43- = AlPO4